Young Company budgets sales of $112,900,000, fixed costs of $25,000,000, and variable costs of $66,611,000. What is the contribution margin ratio for Young Company?
______ %

b. If the contribution margin ratio for Martinez Company is 40%, sales were $34,800,000, and fixed costs were $1,500,000, what was the operating income?
________$
Answer #1

a.Contribution margin=Sales-Variable cost  

=(112,900,000-66,611,000)

=$46289000

Contribution margin ratio=Contribution margin/Sales

=46289000/112,900,000

=41%

b.Contribution margin ratio=Contribution margin/Sales

Contribution margin would be=34,800,000*40%

=$13920000

Less:Fixed cost=1,500,000

Operating income=$12,420,000